### v3

@@ 2.2 Sobolev spaces
Denote $\Omega$ as $[a,b]^{d}$, $D$ as the weak derivative of a single variable function, $D^{\boldsymbol{\alpha}}=D^{\alpha_{1}}_{1}D^{\alpha_{2}}_{2}\ldots D^{\alpha_{d}}_{d}$ as the partial derivative of a multivariable function, where $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\ldots,\alpha_{d}]^{T}$ and $D_{i}$ is the derivative in the $i$-th variable.
Definition 2.1 (Sobolev Spaces [11]).
@@
$W^{n,p}(\Omega):=\left\{f\in L^{p}(\Omega):D^{\boldsymbol{\alpha}}f\in L^{p}(\Omega)\text{ for all }\boldsymbol{\alpha}\in\mathbb{N}^{d}\text{ with }|\boldsymbol{\alpha}|\leq n\right\}$
with a norm $\|f\|_{W^{n,p}(\Omega)}:=\left(\sum_{0\leq|\alpha|\leq n}\left\|D^{\boldsymbol{\alpha}}f\right\|_{L^{p}(\Omega)}^{p}\right)^{1/p}$, if $p<\infty$, and
$\|f\|_{W^{n,\infty}(\Omega)}:=\max_{0\leq|\alpha|\leq n}\left\|D^{\boldsymbol{\alpha}}f\right\|_{L^{\infty}(\Omega)}.$
Furthermore, for $\boldsymbol{f}=(f_{1},f_{2},\ldots,f_{d})$, $\boldsymbol{f}\in W^{1,\infty}(\Omega,\mathbb{R}^{d})$ if and only if $f_{i}\in W^{1,\infty}(\Omega)$ for each $i=1,2,\ldots,d$ and
$\|\boldsymbol{f}\|_{W^{1,\infty}(\Omega,\mathbb{R}^{d})}:=\max_{i=1,\ldots,d}\{\|f_{i}\|_{W^{1,\infty}(\Omega)}\}.$
Definition 2.2 (Sobolev semi-norm [11]).
Let $n\in\mathbb{N}_{+}$ and $1\leq p\leq\infty$. Then we define Sobolev semi-norm $|f|_{W^{n,p}(\Omega)}:=\left(\sum_{|\alpha|=n}\left\|D^{\boldsymbol{\alpha}}f\right\|_{L^{p}(\Omega)}^{p}\right)^{1/p}$, if $p<\infty$, and $|f|_{W^{n,\infty}(\Omega)}:=\max_{|\alpha|=n}\left\|D^{\boldsymbol{\alpha}}f\right\|_{L^{\infty}(\Omega)}$. Furthermore, for $\boldsymbol{f}\in W^{1,\infty}(\Omega,\mathbb{R}^{d})$, we define
$|\boldsymbol{f}|_{W^{1,\infty}(\Omega,\mathbb{R}^{d})}:=\max_{i=1,\ldots,d}\{|f_{i}|_{W^{1,\infty}(\Omega)}\}.$
@@ 3 Deep Super ReLU Networks
Corollary 3.1 ([42]).
For any $f\in W^{n,\infty}((0,1)^{d})$ with $\|f\|_{W^{n,\infty}((0,1)^{d})}\leq 1$, any $N,L,m\in\mathbb{N}_{+}$ with $NL+2^{\left\lfloor\log_{2}N\right\rfloor}\geq\max\{d,n\}$ and $L\geq\left\lceil\log_{2}N\right\rceil$, there is a $\sigma_{2}$-NN $\varphi(\boldsymbol{x})$ with the width $\mathcal{O}(N\log N)$ and depth $\mathcal{O}(L\log L)$ such that
$\|f(\boldsymbol{x})-\varphi(\boldsymbol{x})\|_{W^{m,\infty}((0,1)^{d})}\leq C_{11}(n,d,m)N^{-2(n-m)/d}L^{-2(n-m)/d},$
where $C_{11}$ is the constant independent of $N,L$.
Nevertheless, a limitation of this corollary that uses ReLU and its square as activation functions in DNNs is that the distribution of the square of ReLU is yet to be determined. If the distribution is entirely unknown, training complexity may significantly increase. Hence, this paper introduces a novel DNN architecture that predominantly uses ReLU activation functions, allowing the square of ReLU to appear only in the last several layers. This design allows us to achieve a nearly optimal rate of approximation, reduce the complexity of DNNs as stated in Corollary 3.1, and obtain a better generalization error, making the training process easier. We will rigorously discuss this structure later and show by the numerical experiment that this network performs better in training compared to networks where the activation functions are either all ReLU or all square of ReLU in Sec. 7.
@@
$\mathcal{N}_{C,L}:=$$\{\psi(\boldsymbol{x})=\psi_{2}\circ\boldsymbol{\psi}_{1}(\boldsymbol{x}):\psi_{2}\text{ is a $\sigma_{2}$-NN with depth $L_{2}$,}$
@@ 4 Approximation by DSRNs for Functions in Sobolev Spaces
Theorem 4.1.
For any $f\in W^{n,p}([a,b]^{d})$ for $p\in[1,\infty]$, any $d,n\geq 2$ and $N,L\in\mathbb{N}_{+}$ with $\log N\geq d(\log n+\log d)$, there is a DSRN $\gamma(\boldsymbol{x})$ in $\mathcal{N}_{1,9(L+1)\log_{2}4L}$ with the width $2^{d+6}n^{d+1}(N+d)\log_{2}(8N)$ such that
$\|f(\boldsymbol{x})-\gamma(\boldsymbol{x})\|_{W^{2,p}([a,b]^{d})}\leq 2^{d+7}C_{10}(n,d)\|f(\boldsymbol{x})\|_{W^{n,p}([a,b]^{d})}N^{-2(n-2)/d}L^{-2(n-2)/d},$
where $C_{10}$ is the constant independent of $N,L$.
Remark 4.2.
As we can see, the approximation rate of the DSRN in Theorem 4.1 is
@@
with $\mathcal{O}(N^{2}L)$ parameters, which is better than traditional methods and achieves super-convergence. This rate is achieved based on the bit extraction method, where the super-convergence approximation error decreases faster than the metric entropy of the space of Sobolev functions. Therefore, observing this error order in numerical experiments can be challenging, as discuss the details in [34].
4.1 Proof Sketch for Theorem 4.1
The proof of Theorem 4.1 can be outlined in five parts, and the complete proof is provided in next subsections:
(i): First, define a sequence of subsets of $\Omega$ in Definition 4.5.
Then we define a partition of unity $\{g_{\boldsymbol{m}}\}_{\boldsymbol{m}\in\{1,2\}^{d}}$ on $[a,b]^{d}$ with supp $\text{supp }g_{\boldsymbol{m}}\cap[a,b]^{d}\subset\Omega_{\boldsymbol{m}}$ for each $\boldsymbol{m}\in\{1,2\}^{d}$ in Definition 4.7.
(ii): Next we use the Proposition 4.8 to present $\{s_{\boldsymbol{m}}\}_{\boldsymbol{m}\in\{1,2\}^{d}}$ by $\sigma_{2}$-NNs.
(iii): Third, for each $\Omega_{\boldsymbol{m}}\subset[a,b]^{d}$, where $\boldsymbol{m}\in\{1,2\}^{d}$, we find a function $f_{K,\boldsymbol{m}}$ satisfying
$\|f-f_{K,\boldsymbol{m}}\|_{W^{2,p}(\Omega_{\boldsymbol{m}})}$ $\leq C_{1}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}K^{-(n-2)},$
$\|f-f_{K,\boldsymbol{m}}\|_{W^{1,p}(\Omega_{\boldsymbol{m}})}$ $\leq C_{1}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}K^{-(n-1)},$
(2) $\|f-f_{K,\boldsymbol{m}}\|_{L^{p}(\Omega_{\boldsymbol{m}})}$ $\leq C_{1}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}K^{-n},$
where $C_{1}$ is a constant independent of $K$. Moreover, each $f_{K,\boldsymbol{m}}$ can be expressed as $f_{K,\boldsymbol{m}}=\sum_{|\boldsymbol{\alpha}|\leq n-1}g_{f,\boldsymbol{\alpha},\boldsymbol{m}}(\boldsymbol{x})\boldsymbol{x}^{\boldsymbol{\alpha}}$, where $g_{f,\boldsymbol{\alpha},\boldsymbol{m}}(\boldsymbol{x})$ is a piecewise constant function on $\Omega_{\boldsymbol{m}}$. The proof of this result is based on the Bramble-Hilbert Lemma [4, Lemma 4.3.8].
(iv): The fourth step involves approximating $f_{K,\boldsymbol{m}}$ using DSRN $\gamma_{\boldsymbol{m}}$, following the approach outlined in [26]. This method is suitable for our work because $g_{f,\boldsymbol{\alpha},\boldsymbol{m}}(\boldsymbol{x})$ is a piecewise constant function on $\Omega_{\boldsymbol{m}}$, and the weak derivative of $g_{f,\boldsymbol{\alpha},\boldsymbol{m}}(\boldsymbol{x})$ on $\Omega_{\boldsymbol{m}}$ is zero. This property allows for the use of the $L^{\infty}$ norm approximation method presented in [26]. Thus, we obtain a neural network $\psi_{\boldsymbol{m}}$ with width $\mathcal{O}(N\log N)$ and depth $\mathcal{O}(L\log L)$ such that
(3) $\|f_{K,\boldsymbol{m}}-\gamma_{\boldsymbol{m}}(\boldsymbol{x})\|_{W^{2,p}(\Omega_{\boldsymbol{m}})}$ $\leq C(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}N^{-2(n-2)/d}L^{-2(n-2)/d},$
@@
By combining (iii) and (iv) and setting $K=\lfloor N^{1/d}\rfloor^{2}\lfloor L^{2/d}\rfloor$, we obtain that for each $\boldsymbol{m}\in\{1,2\}^{d}$, there exists a DSRN $\gamma_{\boldsymbol{m}}$ with width $\mathcal{O}(N\log N)$ and depth $\mathcal{O}(L\log L)$ such that
(4) $\|f(\boldsymbol{x})-\gamma_{\boldsymbol{m}}(\boldsymbol{x})\|_{W^{2,p}(\Omega_{\boldsymbol{m}})}$ $\leq C_{10}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}N^{-2(n-2)/d}L^{-2(n-2)/d},$
where $C_{10}$ is a constant independent of $N$ and $L$.
(v): The final step is to combine the sequences $\{s_{\boldsymbol{m}}\}_{\boldsymbol{m}\in\{1,2\}^{d}}$ and $\{\gamma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in\{1,2\}^{d}}$ to construct a network that can approximate $f$ over the entire space $[a,b]^{d}$. We define the sequence $\{s_{\boldsymbol{m}}\}_{\boldsymbol{m}\in\{1,2\}^{d}}$ because $\gamma_{\boldsymbol{m}}$ may not accurately approximate $f$ on $[a,b]^{d}\backslash\Omega_{\boldsymbol{m}}$. The purpose of $s_{\boldsymbol{m}}$ is to remove this portion of the domain and to allow other networks to approximate $f$ on $[a,b]^{d}\backslash\Omega_{\boldsymbol{m}}$.
@@ 4.2 Propositions and Lemmas of $\sigma_{i}$ neural networks
(i) $f(x)=x^{2}$ can be realized exactly by a $\sigma_{2}$ neural network with one hidden layer and two neurons, and $f(x,y)=xy=\frac{(x+y)^{2}-(x-y)^{2}}{4}$ can be realized exactly by a $\sigma_{2}$ neural network with one hidden layer and four neurons.
(ii) Let $P(\boldsymbol{x})$ be a polynomial of total degree $m$ in $\mathbb{R}^{d}$. There is a $\sigma_{2}$-NN of depth $\lceil\log_{2}m\rceil$ and width $2\,(m+1)^{d}$ that exactly represents the polynomial $P(\boldsymbol{x})$.
(iii) For any $s>0$, there is a $\sigma_{2}$-NN with width $2(s+1)$ and depth $\log_{2}m$ that exactly represents $\text{ReLU}^{s+2}$.
Proof 4.4.
(i) is easy to check by $x^{2}=\sigma_{2}(x)+\sigma(-x)$. (ii) is the results in [17]. For (iii), we know that
(5) $\text{ReLU}^{s+2}(x)=x^{s}\cdot\text{ReLU}^{2}(x).$
Then combining (i) and (ii), we have that $\text{ReLU}^{s+2}$ is a $\sigma_{2}$ neural network with width $2(s+1)$ and depth $\log_{2}m$.
@@
Definition 4.5.
Given $K,d\in\mathbb{N}^{+}$, and for any $\boldsymbol{m}=(m_{1},m_{2},\ldots,m_{d})\in\{1,2\}^{d}$, we define $\Omega_{\boldsymbol{m}}:=\prod_{j=1}^{d}\Omega_{m_{j}},$ where
@@
(6) $\Omega_{2}$ $:=\bigcup_{i=0}^{K}\left[\frac{i\cdot(b-a)}{K}-\frac{b-a}{2K},a+\frac{i\cdot(b-a)}{K}+\frac{b-a}{4K}\right]\cap[a,b].$
@@
(7) $s(x):=\begin{cases}2x^{2},~{}&x\in\left[0,\frac{1}{2}\right]\\ -2(x-1)^{2}+1,~{}&x\in\left[\frac{1}{2},1\right]\\ 1,~{}&x\in\left[1,2\right]\\ -2(x-2)^{2}+1,~{}&x\in\left[2,\frac{5}{2}\right]\\ 2(x-3)^{2},~{}&x\in\left[\frac{5}{2},3\right]\\ 0,~{}&\text{otherwise}.\end{cases}$
Figure 2: $s(x)$ in $\mathbb{R}$.
Definition 4.7.
@@
(8) $s_{1}(x)=\sum_{i=0}^{K}s\left(4Kx+1-4i\right),~{}s_{2}(x)=s_{1}\left(x+\frac{1}{2K}\right).$
Then for any $\boldsymbol{m}=(m_{1},m_{2},\ldots,m_{d})\in\{1,2\}^{d}$, we define
(9) $s_{\boldsymbol{m}}(\boldsymbol{x}):=\prod_{j=1}^{d}s_{m_{j}}\left(\frac{x_{j}-a}{b-a}\right)$
for any $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{d})\in\mathbb{R}^{d}$.
Proposition 4.8.
Given $N,L,d\in\mathbb{N}_{+}$ with $\log_{2}N>d\log d$, and the setting $K=\lfloor N^{1/d}\rfloor\lfloor L^{2/d}\rfloor$, $\{s_{\boldsymbol{m}}(\boldsymbol{x})\}_{\boldsymbol{m}\in\{1,2\}^{d}}$ defined in Definition 4.7 satisfies:
(i): $\|s_{\boldsymbol{m}}(\boldsymbol{x})\|_{L^{\infty}([a,b]^{d})}\leq 1$$\|s_{\boldsymbol{m}}(\boldsymbol{x})\|_{W^{1,\infty}([a,b]^{d})}\leq\frac{8K}{b-a}$ and $\|s_{\boldsymbol{m}}(\boldsymbol{x})\|_{W^{2,\infty}([a,b]^{d})}\leq\frac{64K^{2}}{(b-a)^{2}}$ for any $\boldsymbol{m}\in\{1,2\}^{d}$.
(ii): $\{s_{\boldsymbol{m}}(\boldsymbol{x})\}_{\boldsymbol{m}\in\{1,2\}^{d}}$ is a partition of the unity $[a,b]^{d}$ with ${\rm supp}~{}s_{\boldsymbol{m}}(\boldsymbol{x})\cap[a,b]^{d}=\Omega_{\boldsymbol{m}}$ defined in Definition 4.5.
(iii): For any $\boldsymbol{m}\in\{1,2\}^{d}$, we have that there is a $\sigma_{1}$ neural network $\lambda_{\boldsymbol{m},1}(\boldsymbol{x})$ with the width $4N$ and depth $\lceil 2\log_{2}L\rceil$ and a $\sigma_{2}$ neural network $\lambda_{\boldsymbol{m},2}(\boldsymbol{x})$ with the width $6N$ and depth $\lceil 1+\log_{2}d\rceil$, such as
$\lambda_{\boldsymbol{m},2}\circ\lambda_{\boldsymbol{m},1}(\boldsymbol{x})=\prod_{j=1}^{d}s_{m_{j}}\left(\frac{x_{j}-a}{b-a}\right)=s_{\boldsymbol{m}}(\boldsymbol{x}),\boldsymbol{x}\in[a,b]^{d}.$
Proof 4.9.
@@
First, we architect $s(x)$ by a $\sigma_{2}$ neural network. There is a $\sigma_{1}$ neural network $g(x)$ with width 3 and one hidden layer that:
(10) $g(x):=\begin{cases}x,~{}&x\in\left[0,\frac{1}{2}\right]\\ \frac{1}{2},~{}&x\in\Big{[}\frac{1}{2},+\infty\Big{)}\\ 0,~{}&\text{otherwise}.\end{cases}$
Based on (iii) in Lemma 4.3, $g^{2}(x)$ is a $\sigma_{2}$ neural network with width 3 and two hidden layers. Then by direct calculation, we notice that
(11) $s(x)=2[x]_{+}^{2}-4[x-0.5]_{+}^{2}+2[x-1]_{+}^{2}-2[x-2]_{+}^{2}+4[x-2.5]_{+}^{2}-2[x-3]_{+}^{2}.$

### v4

@@ 2.2 Sobolev spaces
Denote $\Omega$ as $a,b]^{d}$, $D$ as the weak derivative of a single variable function, $D^{\bm{\alpha}}=D^{\alpha_{1}}_{1}D^{\alpha_{2}}_{2}\ldots D^{\alpha_{d}}_{d}$ as the partial derivative of a multivariable function, where $\bm{\alpha}=[\alpha_{1},\alpha_{2},\ldots,\alpha_{d}]^{T}$ and $D_{i}$ is the derivative in the $i$-th variable.
Definition 1 (Sobolev Spaces [17]).
@@
$W^{n,p}(\Omega):=\left\{f\in L^{p}(\Omega):D^{\bm{\alpha}}f\in L^{p}(\Omega)\text{ for all }\bm{\alpha}\in\mathbb{N}^{d}\text{ with }|\bm{\alpha}|\leq n\right\}$
with a norm $\|f\|_{W^{n,p}(\Omega)}:=\left(\sum_{0\leq|\alpha|\leq n}\left\|D^{\bm{\alpha}}f\right\|_{L^{p}(\Omega)}^{p}\right)^{1/p}$, if $p<\infty$, and
$\|f\|_{W^{n,\infty}(\Omega)}:=\max_{0\leq|\alpha|\leq n}\left\|D^{\bm{\alpha}}f\right\|_{L^{\infty}(\Omega)}.$
Furthermore, for $\bm{f}=(f_{1},f_{2},\ldots,f_{d})$, $\bm{f}\in W^{1,\infty}(\Omega,\mathbb{R}^{d})$ if and only if $f_{i}\in W^{1,\infty}(\Omega)$ for each $i=1,2,\ldots,d$ and
$\|\bm{f}\|_{W^{1,\infty}(\Omega,\mathbb{R}^{d})}:=\max_{i=1,\ldots,d}\{\|f_{i}\|_{W^{1,\infty}(\Omega)}\}.$
Definition 2 (Sobolev semi-norm [17]).
Let $n\in\mathbb{N}_{+}$ and $1\leq p\leq\infty$. Then we define Sobolev semi-norm $|f|_{W^{n,p}(\Omega)}:=\left(\sum_{|\alpha|=n}\left\|D^{\bm{\alpha}}f\right\|_{L^{p}(\Omega)}^{p}\right)^{1/p}$, if $p<\infty$, and $|f|_{W^{n,\infty}(\Omega)}:=\max_{|\alpha|=n}\left\|D^{\bm{\alpha}}f\right\|_{L^{\infty}(\Omega)}$. Furthermore, for $\bm{f}\in W^{1,\infty}(\Omega,\mathbb{R}^{d})$, we define
$|\bm{f}|_{W^{1,\infty}(\Omega,\mathbb{R}^{d})}:=\max_{i=1,\ldots,d}\{|f_{i}|_{W^{1,\infty}(\Omega)}\}.$
@@ 3 Deep Super ReLU Networks
Corollary 1 ([59]).
Let $f\in W^{n,\infty}((0,1)^{d})$ with $\|f\|_{W^{n,\infty}((0,1)^{d})}\leq 1$. For any $N,L,m\in\mathbb{N}_{+}$ satisfying $NL+2^{\lfloor\log_{2}N\rfloor}\geq\max\{d,n\}$ and $L\geq\lceil\log_{2}N\rceil$, there exists a neural network $\varphi(\bm{x})$ that mixes the ReLU and squared-ReLU activations, whose width is $\mathcal{O}(N\log N)$ and whose depth is $\mathcal{O}(L\log L)$, such that
$\|f-\varphi\|_{W^{m,\infty}((0,1)^{d})}=\mathcal{O}\!\bigl{(}N^{-2(n-m)/d}L^{-2(n-m)/d}\bigr{)}.$
Nevertheless, a limitation of this corollary that uses ReLU and its square as activation functions in DNNs is that the distribution of the square of ReLU is yet to be determined. If the distribution is entirely unknown, training complexity may significantly increase. Hence, this paper introduces a novel DNN architecture that predominantly uses ReLU activation functions, allowing the square of ReLU to appear only in the last several layers. This design allows us to achieve a nearly optimal rate of approximation, reduce the complexity of DNNs as stated in Corollary 1, and obtain a better generalization error, making the training process easier. We will rigorously discuss this structure later and show by the numerical experiment that this network performs better in training compared to networks where the activation functions are either all ReLU or all square of ReLU in Sec. 7.
@@
$\mathcal{N}_{C,L}:=\{\psi(\bm{x})=\psi_{2}\circ\bm{\psi}_{1}(\bm{x}):\psi_{2}\text{ is a $\sigma_{2}$-NN with depth $L_{2}$,}$
@@ 4 Approximation by DSRNs for Functions in Sobolev Spaces
Theorem 2.
For any $f\in W^{n,p}([a,b]^{d})$ for $p\in[1,\infty]$, any $n\geq 2$ and $d,N,L\in\mathbb{N}_{+}$ with $\log N\geq d(\log n+\log d)$, there is a DSRN $\gamma(\bm{x})$ in $\mathcal{N}_{1,9(L+1)\log_{2}4L}$ with the width $2^{d+6}n^{d+1}(N+d)\log_{2}(8N)$ such that
$\|f(\bm{x})-\gamma(\bm{x})\|_{W^{2,p}([a,b]^{d})}\leq 2^{d+7}C_{3}(n,d,a,b)\|f(\bm{x})\|_{W^{n,p}([a,b]^{d})}N^{-2(n-2)/d}L^{-2(n-2)/d},$
where $C_{3}=C_{1}+4^{\frac{2d+p}{2p}}C_{2}$ is the constant independent of $N,L$, and $C_{1},C_{2}$ is defined in Proposition 7.
Remark 1.
As shown in Theorem 2, the approximation rate of the DSRN is
@@
using $\mathcal{O}(N^{2}L)$ parameters, which outperforms traditional methods and achieves super-convergence. This rate is derived via the bit extraction technique, where the approximation error decays faster than the metric entropy of the Sobolev function class. Consequently, observing this rate in numerical experiments can be difficult, as discussed in detail in [52]. Furthermore, the constant $C_{3}$ appearing in Theorem 2, which determines the leading coefficient of the approximation error, depends on the interval endpoints $a$ and $b$. Specifically, $C_{1}$ depends on $\max\{|b-a|^{n-i}\}_{i=0}^{2}$, and $C_{2}$ depends on $\max\{|a|,|b|\}$. Further details are provided in Remark 3.
4.1 Proof Sketch for Theorem 2
The proof of Theorem 2 can be outlined in five parts, and the complete proof is provided in next subsections:
(i): First, define a sequence of subsets of $\Omega$ in Definition 3.
Then we define a partition of unity $\{g_{\bm{m}}\}_{\bm{m}\in\{1,2\}^{d}}$ on $a,b]^{d}$ with $\operatorname{supp}g_{\bm{m}}\cap[a,b]^{d}\subset\Omega_{\bm{m}}$ for each $\bm{m}\in\{1,2\}^{d}$ in Definition 5.
(ii): Next we use the Proposition 4 to present $\{s_{\bm{m}}\}_{\bm{m}\in\{1,2\}^{d}}$ by $\sigma_{2}$-NNs.
(iii): Third, for each $\Omega_{\bm{m}}\subset[a,b]^{d}$, where $\bm{m}\in\{1,2\}^{d}$, we find a function $f_{K,\bm{m}}$ satisfying
$\|f-f_{K,\bm{m}}\|_{W^{2,p}(\Omega_{\bm{m}})}$ $\leq C_{1}(n,d,|b-a|)\|f\|_{W^{n,p}(\Omega_{\bm{m}})}K^{-(n-2)},$
$\|f-f_{K,\bm{m}}\|_{W^{1,p}(\Omega_{\bm{m}})}$ $\leq C_{1}(n,d,|b-a|)\|f\|_{W^{n,p}(\Omega_{\bm{m}})}K^{-(n-1)},$
$\|f-f_{K,\bm{m}}\|_{L^{p}(\Omega_{\bm{m}})}$ $\leq C_{1}(n,d,|b-a|)\|f\|_{W^{n,p}(\Omega_{\bm{m}})}K^{-n},$ (4.1)
where $C_{1}$ is a constant independent of $K$. Moreover, each $f_{K,\bm{m}}$ can be expressed as $f_{K,\bm{m}}=\sum_{|\bm{\alpha}|\leq n-1}g_{f,\bm{\alpha},\bm{m}}(\bm{x})\bm{x}^{\bm{\alpha}}$, where $g_{f,\bm{\alpha},\bm{m}}(\bm{x})$ is a piecewise constant function on $\Omega_{\bm{m}}$. The proof of this result is based on the Bramble-Hilbert Lemma [6, Lemma 4.3.8].
(iv): The fourth step involves approximating $f_{K,\bm{m}}$ using DSRN $\gamma_{\bm{m}}$, following the approach outlined in [39]. This method is suitable for our work because $g_{f,\bm{\alpha},\bm{m}}(\bm{x})$ is a piecewise constant function on $\Omega_{\bm{m}}$, and the weak derivative of $g_{f,\bm{\alpha},\bm{m}}(\bm{x})$ on $\Omega_{\bm{m}}$ is zero. This property allows for the use of the $L^{\infty}$ norm approximation method presented in [39]. Thus, we obtain a neural network $\psi_{\bm{m}}$ with width $\mathcal{O}(N\log N)$ and depth $\mathcal{O}(L\log L)$ such that
$\|f_{K,\bm{m}}-\gamma_{\bm{m}}(\bm{x})\|_{W^{2,p}(\Omega_{\bm{m}})}$ $\leq C(n,d)\|f\|_{W^{n,p}(\Omega_{\bm{m}})}N^{-2(n-2)/d}L^{-2(n-2)/d},$ (4.2)
@@ 4.1 Proof Sketch for Theorem 2
By combining (iii) and (iv) and setting $K=\lfloor N^{1/d}\rfloor^{2}\lfloor L^{2/d}\rfloor$, we obtain that for each $\bm{m}\in\{1,2\}^{d}$, there exists a DSRN $\gamma_{\bm{m}}$ with width $\mathcal{O}(N\log N)$ and depth $\mathcal{O}(L\log L)$ such that
$\|f(\bm{x})-\gamma_{\bm{m}}(\bm{x})\|_{W^{2,p}(\Omega_{\bm{m}})}$ $\leq C_{3}(n,d,a,b)\|f\|_{W^{n,p}(\Omega_{\bm{m}})}N^{-2(n-2)/d}L^{-2(n-2)/d},$ (4.3)
where $C_{3}$ is a constant independent of $N$ and $L$.
(v): The final step is to combine the sequences $\{s_{\bm{m}}\}_{\bm{m}\in\{1,2\}^{d}}$ and $\{\gamma_{\bm{m}}\}_{\bm{m}\in\{1,2\}^{d}}$ to construct a network that can approximate $f$ over the entire space $a,b]^{d}$. We define the sequence $\{s_{\bm{m}}\}_{\bm{m}\in\{1,2\}^{d}}$ because $\gamma_{\bm{m}}$ may not accurately approximate $f$ on $a,b]^{d}\backslash\Omega_{\bm{m}}$. The purpose of $s_{\bm{m}}$ is to remove this portion of the domain and to allow other networks to approximate $f$ on $a,b]^{d}\backslash\Omega_{\bm{m}}$.
@@ 4.2 Propositions and Lemmas of $\sigma_{i}$ neural networks
(i) $f(x)=x^{2}$ can be realized exactly by a $\sigma_{2}$ neural network with one hidden layer and two neurons, and $f(x,y)=xy=\frac{(x+y)^{2}-(x-y)^{2}}{4}$ can be realized exactly by a $\sigma_{2}$ neural network with one hidden layer and four neurons. $x$ can be realized exactly by a $\sigma_{2}$ neural network with one hidden layer and four neurons or $\sigma_{1}$ neural network with one hidden layer and two neurons.
(ii) Let $P(\bm{x})$ be a polynomial of total degree $m$ in $\mathbb{R}^{d}$. There is a $\sigma_{2}$-NN of depth $\lceil\log_{2}m\rceil$ and width $2\,(m+1)^{d}$ that exactly represents the polynomial $P(\bm{x})$.
(iii) For any $s>0$, there is a $\sigma_{2}$-NN with width $2(s+1)$ and depth $\log_{2}m$ that exactly represents $\operatorname{ReLU}^{s+2}$.
Proof.
(i) is easy to check by $x^{2}=\sigma_{2}(x)+\sigma_{2}(-x)$, $\frac{1}{2}((x+1)^{2}-x^{2}-1)=x$ and $x=\sigma_{1}(x)+\sigma_{1}(-x)$. (ii) is the results in [25]. For (iii), we know that
$\operatorname{ReLU}^{s+2}(x)=x^{s}\cdot\operatorname{ReLU}^{2}(x).$ (4.4)
Then combining (i) and (ii), we have that $\operatorname{ReLU}^{s+2}$ is a $\sigma_{2}$ neural network with width $2(s+1)$ and depth $\log_{2}m$. ∎
@@
Definition 3.
Given $K,d\in\mathbb{N}^{+}$, and for any $\bm{m}=(m_{1},m_{2},\ldots,m_{d})\in\{1,2\}^{d}$, we define $\Omega_{\bm{m}}:=\prod_{j=1}^{d}\Omega_{m_{j}},$ where
@@
$\Omega_{2}$ $:=\bigcup_{i=0}^{K}\left[\frac{i\cdot(b-a)}{K}-\frac{b-a}{2K},a+\frac{i\cdot(b-a)}{K}+\frac{b-a}{4K}\right]\cap[a,b].$ (4.5)
@@
$s(x):=\begin{cases}2x^{2},~&x\in\left[0,\frac{1}{2}\right]\\ -2(x-1)^{2}+1,~&x\in\left[\frac{1}{2},1\right]\\ 1,~&x\in\left[1,2\right]\\ -2(x-2)^{2}+1,~&x\in\left[2,\frac{5}{2}\right]\\ 2(x-3)^{2},~&x\in\left[\frac{5}{2},3\right]\\ 0,~&\text{otherwise}.\end{cases}$ (4.6)
Definition 5.
@@
$s_{1}(x)=\sum_{i=0}^{K}s\left(4Kx+1-4i\right),~s_{2}(x)=s_{1}\left(x+\frac{1}{2K}\right).$ (4.7)
Then for any $\bm{m}=(m_{1},m_{2},\ldots,m_{d})\in\{1,2\}^{d}$, we define
$s_{\bm{m}}(\bm{x}):=\prod_{j=1}^{d}s_{m_{j}}\left(\frac{x_{j}-a}{b-a}\right)$ (4.8)
for any $\bm{x}=(x_{1},x_{2},\ldots,x_{d})\in\mathbb{R}^{d}$.
Proposition 4.
Given $N,L,d\in\mathbb{N}_{+}$ with $\log_{2}N>d\log d$, and the setting $K=\lfloor N^{1/d}\rfloor\lfloor L^{2/d}\rfloor$, $\{s_{\bm{m}}(\bm{x})\}_{\bm{m}\in\{1,2\}^{d}}$ defined in Definition 5 satisfies:
(i): $\|s_{\bm{m}}(\bm{x})\|_{L^{\infty}([a,b]^{d})}\leq 1$$\|s_{\bm{m}}(\bm{x})\|_{W^{1,\infty}([a,b]^{d})}\leq\frac{8K}{b-a}$ and $\|s_{\bm{m}}(\bm{x})\|_{W^{2,\infty}([a,b]^{d})}\leq\frac{64K^{2}}{(b-a)^{2}}$ for any $\bm{m}\in\{1,2\}^{d}$.
(ii): $\{s_{\bm{m}}(\bm{x})\}_{\bm{m}\in\{1,2\}^{d}}$ is a partition of the unity $a,b]^{d}$ with ${\rm supp}~s_{\bm{m}}(\bm{x})\cap[a,b]^{d}=\Omega_{\bm{m}}$ defined in Definition 3.
(iii): For any $\bm{m}\in\{1,2\}^{d}$, we have that there is a $\sigma_{1}$ neural network $\lambda_{\bm{m},1}(\bm{x})$ with the width $4N$ and depth $\lceil 2\log_{2}L\rceil$ and a $\sigma_{2}$ neural network $\lambda_{\bm{m},2}(\bm{x})$ with the width $6N$ and depth $\lceil 1+\log_{2}d\rceil$, such as
$\lambda_{\bm{m},2}\circ\lambda_{\bm{m},1}(\bm{x})=\prod_{j=1}^{d}s_{m_{j}}\left(\frac{x_{j}-a}{b-a}\right)=s_{\bm{m}}(\bm{x}),\bm{x}\in[a,b]^{d}.$
Proof.
@@
First, we architect $s(x)$ by a $\sigma_{2}$ neural network. By direct calculation, we notice that
$s(x)=2[x]_{+}^{2}-4[x-0.5]_{+}^{2}+2[x-1]_{+}^{2}-2[x-2]_{+}^{2}+4[x-2.5]_{+}^{2}-2[x-3]_{+}^{2}.$ (4.9)